# Probing Self-Interacting Dark Matter via Gravitational-Wave Background from Eccentric Supermassive Black Hole Mergers

### v1

@@ Appendix A Polynomials for NFW Parameters
Eq. (6) describes the relation between the halo mass and the total stellar mass of galaxies using a double power-law distribution, where the polynomial can be expressed as
@@
where the parameters of the polynomial we select with $B=11.83,\mu=0.18,C=0.047,\nu=-0.4,D=0.728,\eta=-0.16,F=0.052,E=0.92$. For the polynomial in Eq. (8), by fitting the data points from Table. 3, we can get the values with respect to redshift .
Table 3: The relationship between the values of the polynomial in Eq. (8) and the variation in redshift.
@@ Appendix B Orbital Dynamics of the Binary
For a binary system with total mass $M_{\mathrm{tot}}$, component masses $m_{1}$ and $m_{2}$, mass ratio $q=m_{2}/m_{1}$, where $q\in\left(0,1\right)$, the position of the $i$-th black hole in the binary system can be written as [123, 124],
@@
where $a_{i}$ is the corresponding orbital semi-major axis. And the energy and angular momentum of each component are given as
@@
where $M_{i}$ satisfies $M_{1}=\frac{m_{2}^{3}}{M_{\mathrm{tot}}}$ and $M_{2}=\frac{m_{1}^{3}}{M_{\mathrm{tot}}}$, representing the effective central mass governing the orbital motion of each individual star. Then the total energy and total angular momentum can be written as
@@
$\dot{\phi_{i}}$ $=\frac{L_{i}}{m_{i}r^{2}_{i}}=\frac{\sqrt{GM_{i}a_{i}\left(1-e^{2}\right)}}{r^{2}_{i}},$ (50)
@@
where we use the relation $\frac{1}{T}\int_{0}^{T}\mathrm{d}t=\int_{0}^{2\pi}\frac{\mathrm{d}\phi}{2\pi}(1-e^{2})^{\frac{3}{2}}(1+e\mathrm{cos}\phi)^{-2}$ [125]. We note that $N_{1}$ and $N_{2}$ are also functions of position, and their values are determined by Eq. (18) mentioned earlier. Here we do not substitute the specific form of the density profile because for SIDM models, the power-law index is not constant near the transition radius. Therefore, it is necessary to obtain the density profile numerically before calculating the change rates of energy and angular momentum at each position. We sum the contributions of the two components and incorporate the effects of gravitational radiation. Utilizing the Eq. (47) between the semi-major axis, eccentricity, and the energy and angular momentum, we can obtain
@@ Appendix C Number density of black hole mergers
Table 4: Parameter values used in the model.
@@
$m_{\psi z}$ $0.11$ $\beta_{t0}$ $-0.5$
$\alpha_{\psi 0}$ $-1.21$ $\gamma_{t0}$ $-1$
$\alpha_{\psi z}$ $-0.03$
In Eq. (27), we convert the number density of supermassive black hole mergers into the merger event density of galaxies, and the polynomial in Eq. (28) can be expressed as
@@
We have listed the values of the parameters in these polynomials in Table. 4 as [62]. We note that the redshift $z^{\prime}$ used in these polynomials corresponds to the onset of galaxy mergers, and its relationship with the current redshift $z$ can be simply expressed as
$t(z)-t(z^{\prime})=T_{\mathrm{g}-\mathrm{g}}(z^{\prime}),\;t(z=0)=13.79\mathrm{Gyr},$ (61)
where the relationship between redshift and time is described using the standard cosmological model.
@@
Figure 7: The posterior distributions of dark matter model parameters obtained when adopting an eccentricity distribution of $e_{0}=\mathcal{N}[0.7,0.3]$ indicate higher estimates for both the transition velocity and the cross-section constant compared to those in Fig. 4, along with a smaller normalization coefficient $\psi_{0}$.

### v2

@@ Acknowledgements.
DATA AVAILABILITY
The data that support the findings of this article are not publicly available upon publication because it is not technically feasible and/or the cost of preparing, depositing, and hosting the data would be prohibitive within the terms of this research project. The data are available from the authors upon reasonable request.
@@ Appendix A POLYNOMIALS FOR NFW PARAMETERS
Equation (6) describes the relation between the halo mass and the total stellar mass of galaxies using a double power-law distribution, where the polynomial can be expressed as [68]
@@
with the parameters of the polynomial we select being $B=11.83,\mu=0.18,C=0.047,\nu=-0.4,D=0.728,\eta=-0.16,F=0.052,E=0.92$. For the polynomial in Eq. (8), by fitting the data points from Table 3, we can get the values with respect to redshift .
Table 3: The relationship between the values of the polynomial in Eq. (8) and the variation in redshift [124].
@@ Appendix B ORBITAL DYNAMICS OF THE BINARY
For a binary system with total mass $M_{\mathrm{tot}}$, component masses $m_{1}$ and $m_{2}$, and mass ratio $q=m_{2}/m_{1}$, where $q\in\left(0,1\right)$, the position of the $i$-th black hole in the binary system can be written as [125, 126]
@@
where $a_{i}$ is the corresponding orbital semimajor axis. And the energy and angular momentum of each component are given as
@@
where $M_{i}$ satisfies $M_{1}=\frac{m_{2}^{3}}{M_{\mathrm{tot}}^{2}}$ and $M_{2}=\frac{m_{1}^{3}}{M_{\mathrm{tot}}^{2}}$, representing the effective central mass governing the orbital motion of each individual star. Then the total energy and total angular momentum can be written as
@@
$\dot{\phi_{i}}$ $=\frac{L_{i}}{m_{i}r^{2}_{i}}=\frac{\sqrt{GM_{i}a_{i}\left(1-e^{2}\right)}}{r^{2}_{i}}.$ (50)
@@
where we use the relation $\frac{1}{T}\int_{0}^{T}\mathrm{d}t=\int_{0}^{2\pi}\frac{\mathrm{d}\phi}{2\pi}(1-e^{2})^{\frac{3}{2}}(1+e\mathrm{cos}\phi)^{-2}$ [127]. We note that $N_{1}$ and $N_{2}$ are also functions of position, and their values are determined by Eq. (18) mentioned earlier. Here we do not substitute the specific form of the density profile because for SIDM models, the power-law index is not constant near the transition radius. Therefore, it is necessary to obtain the density profile numerically before calculating the change rates of energy and angular momentum at each position. We sum the contributions of the two components and incorporate the effects of gravitational radiation. Utilizing Eq. (47) between the semimajor axis, eccentricity, and the energy and angular momentum, we can obtain
@@ Appendix C NUMBER DENSITY OF BLACK HOLE MERGERS
Table 4: Parameter values used in the model as [63].
@@
$m_{\psi z}$ $0.11$ $\beta_{t0}$ 0.5-0.5
$\alpha_{\psi 0}$ 1.21-1.21 $\gamma_{t0}$ 1-1
$\alpha_{\psi z}$ 0.03-0.03
In Eq. (27), we convert the number density of supermassive black hole mergers into the merger event density of galaxies, and the polynomial in Eq. (28) can be expressed as [63]
@@
We have listed the values of the parameters in these polynomials in Table 4 as [63]. We note that the redshift $z^{\prime}$ used in these polynomials corresponds to the onset of galaxy mergers, and its relationship with the current redshift $z$ can be simply expressed as
$t(z)-t(z^{\prime})=T_{\mathrm{g}-\mathrm{g}}(z^{\prime}),\;t(z=0)=13.79~\mathrm{Gyr},$ (61)
where the relationship between redshift and time is described using the standard cosmological model,
@@
Figure 7: The posterior distributions of dark matter model parameters obtained when adopting an eccentricity distribution of $e_{0}=\mathcal{N}[0.7,0.3]$ indicate higher estimates for both the transition velocity and the cross section constant compared to those in Fig. 4, along with a smaller normalization coefficient $\psi_{0}$.